Jan is twice as old as Phyllis. In ten years Jan will be three times as old as Phyllis. How old are both Jan and Phyllis now?
Answer by AnlytcPhil(1806)


Jan is twice as old as Phyllis. In ten years Jan will be three times as old as
Phyllis. How old are both Jan and Phyllis now?

This problem is botched.  It's impossible because if Jan is twice as old as
Phyllis is now, then in ten years Jan will be LESS THAN TWICE as old as
Phyllis will be then, not MORE THAN TWICE.  So she couldn't possibly be three
times as old in ten years.  

I will make a new feasible problem by swapping the "twice" and the "three
times":

Jan is three times as old as Phyllis. In ten years Jan will be twice as old as
Phyllis. How old are both Jan and Phyllis now?

Jan is twice as old as Phyllis.

                         J = 3P

In ten years Jan will be twice as old as Phyllis. 

In ten years Jan will be J+10
In ten years Phyllis will be P+10

J+10 will be 2 times P+10

                      J+10 = 2(P+10)

So you have the system of equations

                         J = 3P

                      J+10 = 2(P+10)

Solve that system, get P = 10, J = 30
  
Edwin
AnlytcPhil@aol.com